# The role of synthetic data in Multilingual, Multi-cultural AI systems: Lessons from Indic Languages

### v1

@@ 8 Ethical Considerations
Our discussion of ethical considerations is guided by the framework proposed by Bender & Friedman (2018).
@@
Annotators were recruited through a professional external services company. All annotators assigned to a given data point were native speakers of the language represented in the data. Table 7 summarizes the annotator demographics (education, region, age distribution, and gender). Each data worker was compensated at the rate of \$2 per data point.
@@
Table 7: Participant demographics summary
Appendix A Appendix
A.1 Rubrics used for the quality evaluation of the synthetic data
Table 8: Evaluation Rubrics for AI Tasks
Task Rubric Description
@@ A.1 Rubrics used for the quality evaluation of the synthetic data
Creative Writing Instruction adherence Assesses if the output strictly follows all constraints and guidelines provided in the prompt.
@@
Fluency Evaluates the naturalness, grammatical correctness, and readability of the generated text.
@@
Narrative coherence Checks for logical consistency in the plot, character development, and thematic elements.
@@
Reasoning Tasks Answer adequacy Determines if the final answer is correct, complete, and directly addresses the core question.
@@
Context adherence Measures whether the reasoning remains faithful to the provided context, avoiding external facts.
@@
Instruction adherence Verifies that the output’s structure, format, and steps match the user’s instructions.
@@
Fluency and readability Assesses the clarity, logical flow, and ease of understanding of the explanation.
@@
Problematic content and cultural relevance Scrutinizes the response for harmful stereotypes and ensures it is culturally sensitive and appropriate.
@@
Multi-turn Dialog Persona Adherence Evaluates the model’s ability to consistently maintain a specific character or role throughout the conversation.
@@
Topic adherence Checks if the conversation remains focused on the established topic or transitions logically.
@@
Linguistic plausibility Assesses whether the dialogue sounds natural, human-like, and contextually appropriate.
@@
Repetitiveness Measures the degree to which the model avoids unnecessarily repeating phrases or ideas.
@@
Toxicity check Ensures the response is free from any offensive, harmful, or inappropriate content.
@@
Instruction adherence Verifies that the model follows meta-instructions given by the user during the dialogue.
@@
Summarization Coverage Determines if the summary successfully captures all essential points from the source text.
@@
Factual accuracy Checks that the summary correctly represents the information and facts from the original document.
@@
Conciseness Evaluates whether the summary is significantly shorter than the source while retaining critical information.
@@
Coherence and logical flow Assesses if the summary is well-structured, logically organized, and easy to follow.
@@
Style and tone Measures how well the summary reflects the style and tone of the original text.
@@
Translation Semantic correctness Assesses whether the meaning, intent, and nuance of the source text are accurately conveyed.
@@
Fluency correctness Evaluates the grammatical accuracy and naturalness of the translated text in the target language.
@@
Domain appropriateness Checks if the terminology is correct and suitable for the specific subject matter (e.g., legal, medical).
@@
Style and tone Determines if the translation successfully captures the original author’s writing style and emotional tone.
@@
Completeness Verifies that the entire source text has been translated without any omissions or additions.
A.2 Results from the quality evaluations of the synthetic data
@@
Figures  3 and 4 show the distribution of the scores received for the tasks we evaluated across the languages for the tasks. Expert human evaluators have consistently given a score of 2 across languages and tasks, indicating the high quality of Updesh. The disagreements between humans and the LLM is however unclear from these plots. We thereby, performed a thorough inter annotator analysis, the details of which are present in the next figures.
@@
Figure 6: Agreement between human and LLM evaluators per task and language respectively
Continuing the claims made from Figure 1, Figure 5 provides a clearer view of the tasks on which humans and LLMs are likely to agree or disagree. We find that the largest disagreements occur in tasks such as assessing the linguistic plausibility of a given text in a regional language. Furthermore, LLMs struggle with evaluating long-context tasks, such as evaluating whether the same persona is maintained throughout a multi-turn conversation in a regional language. There is also a notable divergence between what human evaluators consider fluent in a relatively low-resource language and what an LLM deems fluent. In contrast, we observe considerable agreement in tasks like toxicity detection and problematic content flagging. LLMs also perform reasonably well at identifying whether a text is culturally relevant, but issues arise when the evaluation requires more fine-grained judgments of multilinguality and multiculturalism. We do not identify any specific trends language or task-wise as apparent in Figure6.
A.3 Data Quality Assesment of Reasoning Data
Language analytical brain fermi fs (cot_flow) math mcq rc text_class
@@ A.3 Data Quality Assesment of Reasoning Data
A.4 Hyperparameters
Table 10: Training hyperparameters used for all our experiments
Hyperparameter Value
Base Model phi4-base-hf / llama-3.1-8b-hf
Sequence Length 65,536
Effective Batch Size 8192
Number of Epochs 3
Optimizer AdamW Torch Fused
Learning Rate 1.0e-05
LR Scheduler Cosine
Adam Beta1 0.9
Adam Beta2 0.95
Max Grad Norm 1.0
Warmup Ratio 0.03
Weight Decay 0.1
NEFTune Noise Alpha 5
Precision BF16
Flash Attention True
Gradient Checkpointing True
A.5 Evaluation Datasets
Table 11: Evaluation datasets
Dataset Source
NLU Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021)
MMLU Indic (MMLU-I) SarvamAI
ARC Indic (ARC-I) SarvamAI
BoolQ Indic (BoolQ-I) SarvamAI
TriviaQA Indic (TVQA-I) SarvamAI
BeleBele (Bele) Bandarkar et al. (2024)
INCLUDE (INCL) Romanou et al. (2025)
Global MMLU (GMMLU) Singh et al. (2025)
NLG Extreme Summarization (Xsum) Singh et al. (2024a)
Flores English to Other (Flores EnXX) Goyal et al. (2022)
Flores XX-En (Flores XXEn)
IN22-Conv (in22-conv-doc) Gala et al. (2023); Gumma et al. (2025)
IF Multilingual Instruction Following Evaluation (mIFEval) Zhou et al. (2023b)
Multilingual Instruction Following Benchmark (mIFBench) Pyatkin et al. (2025)
A.6 Prompt used for conversion of IFEval and IFBench
Translation Generation Prompt INSTRUCTION: You are a professional translator. Your sole task is to translate the provided English text into {target_language}. Rules: Output only the translation—nothing else (no explanations, no commentary). Do not answer or solve any task contained in the text; just translate it. Preserve formatting, punctuation, numbers, placeholders, and bracketed tokens exactly as they appear. Translate idiomatically and naturally for a fluent native speaker of {target_language}. If the input is an instruction (e.g., “Write a resume for …”), translate the instruction itself—do not perform it. Do not introduce extra formatting (e.g., don’t create JSON blocks unless the source already has them). SOURCE LANGUAGE: English TARGET LANGUAGE:
{target_language}
TEXT TO TRANSLATE:
{source_text}
OUTPUT REQUIREMENT:
Return only the translated text in {target_language} with formatting faithfully preserved.
A.7 Sample Prompts Used for Evaluation

### v2

@@ 8 Ethical Considerations
Our discussion of ethical considerations is guided by the framework proposed by Bender and Friedman (2018).
@@
Annotators were recruited through a professional external services company. All annotators assigned to a given data point were native speakers of the language represented in the data. Table 6 summarizes the annotator demographics (education, region, age distribution, and gender). Each data worker was compensated at the rate of \$2 per data point, which is significantly higher than the average for an regular annotation task.
@@
Table 6: Participant demographics summary.
Reproducibility
We provide a detailed reproducibility statement in Appendix A.
Appendix A Reproducibility Statement
To ensure full reproducibility of our pipeline—spanning data generation, model fine-tuning, downstream evaluation, and synthetic data quality assessment — we provide our complete codebase and sample generated data in the supplementary material. The submission zip contains organized data and code folders, with the code structured into seven components corresponding to distinct pipeline stages:
  1. 1.
    selective_translation – Scalable codebase for generating reasoning data through translation from OrcaAgent-Instruct.
  2. 2.
    wiki_bt – Wikipedia-based grounded synthetic data curation pipeline.
  3. 3.
    DCAD-2000 – Modified repository for heuristic-based filtering of standard and multi-turn conversational data.
  4. 4.
    quality_evals – Synthetic data quality evaluation code, including outputs, human annotations, and analysis.
  5. 5.
    model_trainingAxolotl-based fine-tuning scripts for the generated datasets.
  6. 6.
    lm_evaluation_harness – Custom fork integrating additional benchmarks for downstream evaluation.
  7. 7.
    plotting – Analysis and visualization code for experimental results.
Comprehensive instructions are provided in the main README and component-specific README files for straightforward reproduction. Model training configurations with exact hyperparameters are detailed in Appendix 9.
We will publicly release all the following artifacts (code and datasets) under a permissive license.
  • Complete codebase including selective translation pipelines, Wikipedia-grounded synthetic data generation, data filtering methods, model training scripts, and the up-to-date evaluation frameworks
  • Training datasets comprising Updesh Reasoning and Generative subsets across multiple languages.
  • Evaluation datasets featuring GPT-4o translated variants of standard benchmarks (IFEval, IFBench) for reproducibility and future benchmarking.
  • Raw evaluation scores for all models across every dataset, providing complete experimental transparency
  • Human-annotations of synthetic data specifically useful for calibrating / meta-evaluation of LLM evaluators on Indian languages
Figure 6: Framework for Multilingual & Multicultural Synthetic Data Generation
Appendix B Appendix
B.1 Design Considerations
Our framework (Figure˜6) addresses synthetic data generation for multilingual and multicultural contexts throughout the AI lifecycle — pre-training, SFT, RLHF, and evaluation (Viswanathan et al., 2025). Below, we describe the key factors to consider when generating this type of data. While our focus is on SFT data, these design considerations can generalize to other synthetic data types.
Base model capability & Seed data selection
Select foundation models based on performance in target languages on multilingual benchmarks. When language-specific benchmarks are unavailable, use related languages or overall multilingual performance as proxies. Other critical aspects to consider include licensing, cost, and model availability (open-weights vs. restricted). For seed data selection, cover diverse sources and tasks, prioritizing tasks containing cultural knowledge, norms, and values relevant to specific regional contexts.
Data generation strategy
Three primary approaches: (i) Translation from English SFT datasets to transfer critical skills, though risking translationese artifacts (Zhang and Toral, 2019; Vanmassenhove et al., 2021); (ii) Back-translation—using existing unlabeled multilingual datasets through instruction backtranslation (Li et al., 2024) or back-and-forth translation (Nguyen et al., 2024); (iii) Retrieval-augmented generation —leveraging curated native speaker-authored content from web corpora to capture cultural knowledge and linguistic nuances. Translation-based approaches yield weaker correlations with human judgements than language-specific benchmarks (Kreutzer et al., 2025; Wu et al., 2025). Bottom-up approaches grounded in web corpora show superior performance (Shaham et al., 2024; Khan et al., 2024; Doshi et al., 2024) but remain underexplored in multilingual contexts.
Quality metrics
Essential dimensions for assessment: Language correctness—proper language, register, dialect identification (Marchisio et al., 2024); Linguistic acceptability—native speaker fluency and naturalness (Hada et al., 2024b, a); Cultural appropriateness—accurate cultural references, values, and norms; Bias and safety—absence of stereotypes and culturally inappropriate content (Pawar et al., 2025).
Downstream evaluation
Select benchmarks covering all target languages, avoiding English-translated datasets. Include diverse tasks testing cultural knowledge and values. Address benchmark contamination risks (Ahuja et al., 2024) and develop new benchmarks when necessary.
Native speaker involvement
Engage native speakers in seed selection and evaluation. Ensure informed consent addressing cultural considerations and data sovereignty per local regulations. Exclude personally identifiable information from all sources.
B.2 Task Descriptions
Task Descriptions for all the subtasks in Updesh can be found in Table 7 and 8.
Task Type Description
Brain Teasers Puzzles stimulating lateral thinking and creativity
Text Classification Categorization tasks for predefined labels
Reading Comprehension Questions based on understanding and interpreting textual passages
Math Grade-school arithmetic, algebra, and geometry word problems
Table 7: Reasoning task categories and their descriptions.
Task Type Synthesis Method Phases Qwen3-Mode
B.3 Rubrics used for the quality evaluation of the synthetic data
Creative Writing
@@ B.3 Rubrics used for the quality evaluation of the synthetic data
  • Instruction adherence: Assesses if the output strictly follows all constraints and guidelines provided in the prompt.
@@
  • Fluency: Evaluates the naturalness, grammatical correctness, and readability of the generated text.
@@
  • Narrative coherence: Checks for logical consistency in the plot, character development, and thematic elements.
Reasoning Tasks
@@
  • Answer adequacy: Determines if the final answer is correct, complete, and directly addresses the core question.
@@
  • Context adherence: Measures whether the reasoning remains faithful to the provided context, avoiding external facts.
@@
  • Instruction adherence: Verifies that the output’s structure, format, and steps match the user’s instructions.
@@
  • Fluency and readability: Assesses the clarity, logical flow, and ease of understanding of the explanation.
@@
  • Problematic content and cultural relevance: Scrutinizes the response for harmful stereotypes and ensures it is culturally sensitive and appropriate.
Multi-turn Dialog
@@
  • Persona adherence: Evaluates the model’s ability to consistently maintain a specific character or role throughout the conversation.
@@
  • Topic adherence: Checks if the conversation remains focused on the established topic or transitions logically.
@@
  • Linguistic plausibility: Assesses whether the dialogue sounds natural, human-like, and contextually appropriate.
@@
  • Repetitiveness: Measures the degree to which the model avoids unnecessarily repeating phrases or ideas.
@@
  • Toxicity check: Ensures the response is free from any offensive, harmful, or inappropriate content.
@@
  • Instruction adherence: Verifies that the model follows meta-instructions given by the user during the dialogue.
Summarization
@@
  • Coverage: Determines if the summary successfully captures all essential points from the source text.
@@
  • Factual accuracy: Checks that the summary correctly represents the information and facts from the original document.
@@
  • Conciseness: Evaluates whether the summary is significantly shorter than the source while retaining critical information.
@@
  • Coherence and logical flow: Assesses if the summary is well-structured, logically organized, and easy to follow.
@@
  • Style and tone: Measures how well the summary reflects the style and tone of the original text.
Translation
@@
  • Semantic correctness: Assesses whether the meaning, intent, and nuance of the source text are accurately conveyed.
@@
  • Fluency correctness: Evaluates the grammatical accuracy and naturalness of the translated text in the target language.
@@
  • Domain appropriateness: Checks if the terminology is correct and suitable for the specific subject matter (e.g., legal, medical).
@@
  • Style and tone: Determines if the translation successfully captures the original author’s writing style and emotional tone.
@@
  • Completeness: Verifies that the entire source text has been translated without any omissions or additions.
B.4 Results from the quality evaluations of the synthetic data
Figures  7 and 8 show the distribution of the scores received for the tasks we evaluated across the languages for the tasks. Expert human evaluators have consistently given a score of 2 across languages and tasks, indicating the high quality of Updesh. The disagreements between humans and the LLM is however unclear from these plots. We thereby, performed a thorough inter annotator analysis, the details of which are present in the next figures.
Continuing the claims made from Figure 2, Figure 9 provides a clearer view of the tasks on which humans and LLMs are likely to agree or disagree. We find that the largest disagreements occur in tasks such as assessing the linguistic plausibility of a given text in a regional language. Furthermore, LLMs struggle with evaluating long-context tasks, such as evaluating whether the same persona is maintained throughout a multi-turn conversation in a regional language. There is also a notable divergence between what human evaluators consider fluent in a relatively low-resource language and what an LLM deems fluent. In contrast, we observe considerable agreement in tasks like toxicity detection and problematic content flagging. LLMs also perform reasonably well at identifying whether a text is culturally relevant, but issues arise when the evaluation requires more fine-grained judgments of multilinguality and multiculturalism. We do not identify any specific trends language or task-wise as apparent in Figure˜10.
@@ B.4 Results from the quality evaluations of the synthetic data
Figure 10: Agreement between human and LLM evaluators per task and language, respectively
B.5 Generation Hyperparameters
For data synthesis, decoding is performed using nucleus sampling with top_p=0.95top\_p=0.95 and $temperature=1.0$.
B.6 Training Hyperparameters
Hyperparameters for all our training runs could be found in Table 9
Hyperparameter Value
Base Model phi4-base / llama-3.1-8b
Sequence Length 65,536
Effective Batch Size 8192
Number of Epochs 3
Optimizer AdamW
Learning Rate 1e-5
LR Scheduler Cosine
Adam Betas (0.9, 0.95)
Max Grad Norm 1.0
Warmup Ratio 0.03
Weight Decay 0.1
NEFTune Noise Alpha Jain et al. (2024) 5
Precision BF16
Flash Attention True
Gradient Checkpointing True
Table 9: Training hyperparameters used for all our experiments.
B.7 Data Quality Assesment of Reasoning Data
Language Analytical Brain Fermi Fs-CoT Math MCQ RC Text Class
@@ B.7 Data Quality Assesment of Reasoning Data
B.8 Detailed Results
Model Setting NLU Avg MMLU-I MILU ARC-I BoolQ-I TVQA-I Bele INCL GMMLU
Llama 8B Base 29.33 22.97 25.88 25.26 62.04 26.55 24.06 24.88 23.02
Bactrian-X 45.37 36.80 39.41 26.49 71.21 66.30 41.45 39.14 42.14
IndicAlign 42.14 34.47 39.67 27.35 71.14 59.35 32.09 34.37 38.67
Aya-Collection 42.73 32.69 36.98 28.11 76.24 50.65 46.54 33.84 36.81
Updesh 42.89 34.40 36.13 29.67 77.51 46.13 51.30 30.56 37.43
Phi-4 Base 53.17 43.38 49.22 29.99 79.56 68.25 54.55 45.46 54.97
Bactrian-X 49.25 39.91 45.09 29.40 74.40 67.30 48.32 41.72 47.89
IndicAlign 51.71 44.99 48.25 32.23 76.26 61.68 51.68 44.71 53.92
Aya-Collection 54.41 45.08 50.57 33.01 72.62 71.80 60.71 48.02 53.47
Updesh 56.61 48.98 51.19 32.08 81.38 60.14 74.42 47.24 57.47
Table 11: Performance comparison of Llama 8B and Phi-4 variants across Indic NLU Tasks. All entries are accuracy (higher is better)
Model Setting NLG Avg Flores En-XX Flores XX-En XSum IN22-Conv-Doc En-XX IN22-Conv-Doc XX-En
Llama 8B Base 8.91 1.45 41.71 0.16 0.60 0.60
Bactrian-X 23.81 28.85 50.98 0.21 19.51 19.51
IndicAlign 15.67 32.12 3.20 12.84 15.09 15.09
Aya-Collection 9.69 28.38 0.46 0.23 9.69 9.69
Updesh 36.41 44.00 51.88 25.54 30.31 30.31
Phi4 Base 34.20 30.23 56.57 17.59 33.31 33.31
Bactrian-X 24.45 26.78 51.86 0.31 21.60 21.71
IndicAlign 13.77 32.13 0.59 0.28 17.89 17.94
Aya-Collection 13.93 30.16 1.58 0.37 18.77 18.78
Updesh 36.56 45.82 59.55 21.66 27.81 27.94
Table 12: Performance comparison of Llama 8B and Phi-4 variants across Indic NLG. All entries are ChrF scores (higher is better).
Model Variant Avg as bn en gu hi kn ml mr ne or pa ta te ur
Llama 8B Base ZS 28.98 22.22 29.61 36.08 29.97 29.19 30.41 30.82 30.72 23.02 31.14 30.14 29.91 28.97 23.44
Bactrian 45.65 38.22 45.73 68.19 44.06 50.88 43.41 44.07 48.05 40.12 40.24 44.90 40.81 40.35 50.11
IndicAlign 41.03 32.11 42.04 63.79 42.05 44.55 42.36 41.30 43.38 31.84 38.87 41.52 40.79 39.83 30.00
Aya 44.30 42.78 42.43 63.04 42.07 45.66 43.28 43.76 43.15 36.03 38.52 43.06 41.74 40.85 53.89
Updesh 45.23 46.44 41.97 59.84 44.52 42.35 44.95 45.56 42.99 35.63 43.99 44.08 41.82 41.98 57.11
Phi-4 Base ZS 53.36 42.22 52.77 75.64 53.93 60.37 51.55 49.40 53.25 48.14 47.51 54.21 48.00 47.84 62.22
Bactrian 50.16 40.67 49.39 77.47 49.21 56.36 45.57 45.53 50.54 45.06 44.47 50.60 41.87 41.80 63.67
IndicAlign 52.86 49.33 51.13 75.84 50.74 55.49 50.33 50.88 49.38 49.91 42.72 51.22 48.07 47.44 67.56
Aya 55.80 52.89 53.41 76.41 55.16 60.23 54.50 51.63 55.01 52.87 46.25 55.46 50.10 49.33 68.00
Updesh 59.71 70.33 55.77 75.45 56.91 61.82 56.43 54.20 56.81 58.21 53.35 57.59 51.49 52.04 75.56
Table 13: NLU Performance: Evaluation across Indic languages including English. Language codes: as: Assamese, bn: Bengali, en: English, gu: Gujarati, hi: Hindi, kn: Kannada, ml: Malayalam, mr: Marathi, ne: Nepali, or: Odia, pa: Punjabi, ta: Tamil, te: Telugu, ur: Urdu.
Model Variant Avg as bn gu hi kn ml mr ne or pa ta te ur
Llama 8B Base ZS 9.08 7.05 7.87 7.76 11.05 9.22 9.04 9.77 11.83 8.48 10.26 7.30 9.24 9.22
Bactrian 24.24 20.32 29.79 24.24 33.13 22.87 23.97 19.58 28.30 18.26 21.63 24.24 23.32 25.51
IndicAlign 15.80 13.09 18.29 12.27 26.29 15.34 14.75 17.43 21.57 6.65 13.71 17.22 14.66 14.14
Aya 9.87 7.49 12.98 10.83 8.62 11.07 12.74 8.04 11.86 6.42 8.34 12.00 8.78 9.10
Updesh 36.68 33.80 35.98 34.90 36.81 35.73 37.75 34.15 43.13 32.87 35.60 39.08 39.64 37.42
Phi-4 Base ZS 34.56 25.39 39.47 38.83 46.31 30.02 30.97 32.36 40.76 27.36 40.32 30.73 34.61 32.10
Bactrian 24.96 20.74 30.75 24.90 40.36 22.72 23.68 23.53 33.54 13.29 17.42 25.19 25.27 23.14
IndicAlign 14.08 12.81 16.57 15.35 25.27 12.46 11.59 13.06 20.42 6.95 10.81 15.58 19.52 2.60
Aya 14.29 7.67 21.88 15.44 25.36 5.15 12.57 12.80 23.35 1.91 16.39 12.33 15.49 15.40
Updesh 36.87 32.60 37.66 38.50 40.38 34.21 36.50 35.12 42.16 33.56 36.29 39.99 37.61 34.74
Table 14: NLG Performance: Evaluation across Indic languages (excluding English). Language codes: as: Assamese, bn: Bengali, gu: Gujarati, hi: Hindi, kn: Kannada, ml: Malayalam, mr: Marathi, ne: Nepali, or: Odia, pa: Punjabi, ta: Tamil, te: Telugu, ur: Urdu.
Model Variant Avg awa bho brx hne gbm bgc gom mai mup mni mwr ps hoj sa sat bo
Llama 8B Base ZS 10.13 15.86 14.82 2.64 13.06 13.72 15.41 7.24 9.34 13.66 1.00 15.13 13.29 13.88 5.81 2.80 4.39
Bactrian 19.59 27.70 19.28 4.45 28.41 27.56 26.20 13.23 23.72 26.71 5.00 29.27 20.09 26.89 17.38 4.25 13.29
IndicAlign 9.19 12.45 11.03 0.32 15.52 11.10 15.35 6.00 10.16 13.79 2.23 14.36 7.14 14.43 11.59 0.21 1.38
Aya 6.10 6.97 7.63 0.24 7.28 8.87 9.23 4.28 7.16 8.93 2.16 9.36 3.16 9.78 6.55 0.23 5.82
Updesh 24.40 33.67 32.33 4.30 37.34 30.20 35.61 23.53 26.23 33.13 6.60 36.50 21.46 33.69 21.25 4.55 10.04
Phi-4 Base ZS 22.69 35.45 32.75 5.01 30.05 31.04 26.78 15.74 31.99 27.04 5.73 29.62 25.09 24.09 26.08 4.99 11.61
Bactrian 22.19 31.71 27.96 4.80 31.14 30.38 31.70 15.38 26.43 30.58 5.90 31.13 19.33 30.11 21.13 3.80 13.64
IndicAlign 9.41 12.51 10.26 1.33 12.83 12.65 12.83 9.09 17.01 13.51 1.28 13.52 6.76 10.14 11.03 2.21 3.58
Aya 12.20 17.67 13.22 1.12 16.44 16.89 17.25 7.98 18.41 14.90 0.55 16.13 12.47 14.14 14.86 4.25 8.95
Updesh 23.51 34.32 30.71 4.90 35.04 32.39 28.16 21.88 26.82 31.22 7.29 33.82 21.04 29.12 22.06 4.19 13.15
Table 15: Detailed performance breakdown for out-of-training languages. Language codes: awa: Awadhi, bho: Bhojpuri, brx: Bodo, hne: Chhattisgarhi, gbm: Garhwali, bgc: Haryanvi, gom: Konkani, mai: Maithili, mup: Malvi, mni: Manipuri, mwr: Marwari, ps: Pashto, hoj: Rajasthani, sa: Sanskrit, sat: Santali, bo: Tibetan.
Figure 11: Scatter plots correlating ELO Scores with NLU Average (left), NLG Average (center), and Dataset Wins (right). The Updesh-32K model (UB 32K) consistently outperforms baselines, appearing in the top-right quadrant across all metrics.
Appendix C Cultural Evaluation Framework
To assess the local cultural and community-specific knowledge of Large Language Models (LLMs) within the Indian context, we collected a set of questions in collaboration with a third-party non-profit organization. The dataset consists of 4,399 unique queries that reflect authentic information-seeking behaviors of local populations. These queries span 11 Indian languages (see Figure 12) and cover four high-impact domains: Healthcare, Education, Finance, and Legal. To ensure a robust assessment, we performed a comparative evaluation of all model checkpoints, totaling 91,982 pairwise comparisons (battles). We refer the reader to Hamna et al. (2025) for the collection process, which we replicated.
Domain Themes.
The dataset covers a wide spectrum of community needs and themes, some of which were:
  • Education: Teaching and learning support, and career guidance.
  • Finance: Insurance, savings, and budgeting strategies.
  • Healthcare: Senior care protocols and general wellness habits.
  • Legal: Product/service disputes and family or marriage-related inquiries.
Figure 12: Cultural evaluation language wise distribution
Figure 13: Exanples of a Punjabi and Malayalam query we collected
C.1 Sample Prompts Used for Evaluation